# How to Synthesize Text Data without Model Collapse?

### v1

@@ 2.2 Why does synthetic data fail in language model pre-training?
Finding IV: Distribution shifting cannot be mitigated through data selection. Inspired by recent data selection works (Xie et al., 2023; Albalak et al., 2024), we try to leverage the human-produced data features as a reference distribution for selecting synthetic samples. We apply importance sampling from DSIR (Xie et al., 2023) to filter synthetic data. As shown in Figure 4A, the sampled data still fails to align with human-produced data in the embedding space, even at the boundary regions of the synthetic data. As illustrated in Figure 4B, the training results of selected synthetic samples still fluctuates around the original performance of the synthetic data, indicating that even biased sampling cannot correct the distributional shift.
2.3 Proposed Strategy
Following these lessons so far, due to the coverage collapse and feature over-concentration properties of synthetic data, the best approach is to rely entirely on human-produced data and avoid including synthetic data. However, we are still wondering how synthetic data can be used to enhance human-produced data. This leads to a general guideline for synthetic data: relying solely on synthetic data leads to model collapse, so preserving the primary human-produced data distribution is essential. In that case, we propose token-level editing, which leverages a prior distribution to adjust the data. Our method can maintain the source distribution while improving the source data, called semi-synthetic data.
@@ 3 Token-Level Editing
In this section, we introduce token-level editing as a method for generating semi-synthetic data. Furthermore, we present a theoretical analysis and proof demonstrating that the test squared error of our method has a finite upper bound, regardless of the number of iterations. Consequently, our method not only prevents model collapse but also achieves improved performance.
@@ 3.1 Method
$P(y_{1},\dots,y_{n}\mid x_{1},\dots,x_{t})=\prod_{i=1}^{n}P(y_{i}\mid y_{1},\dots,y_{i-1},x_{1},\dots,x_{t}).$ (1)
@@
Inspired by previous studies of data selection (Mindermann et al., 2022; Ankner et al., 2024; Lin et al., 2024), prior distributions can serve as pointers indicating useless or learnable samples. In this case, we use a pre-trained language model to infer the pretraining corpus. As illustrated in Figure 6, even a model pre-trained on trillions of tokens can not fit the pretraining corpus perfectly. Specifically, $75\%$ is under 0.6 probability. The tokens with both high and low probabilities are the most concentrated, suggesting the potential for data filtering. We leverage this U-shape distribution as a pointer to resample tokens. Specifically, we use a language model as prior distribution to compute each token’s conditional probability $P(\cdot|\textbf{x})$ if the probability exceeds a certain threshold $P(\cdot|\textbf{x})\geq p$, it indicates that the token is easy to learn, and we proceed with resampling at that point.
@@ 3.2 Theoretical Analysis
To gain deeper mathematical insights, we utilize an analytical framework of the linear model and adopt notations in prior research (Mobahi et al., 2020; Dohmatob et al., 2024a; Gerstgrasser et al., 2024). This theoretical framework primarily considers a linear model that iteratively trains on its own generated data, similar to pipelines like self-play and self-distillation, but without complex constraints. The process involves training continuously on the data generated by the previous generation of the model. Dohmatob et al. (2024a) point out that with iterative training, test errors accumulate progressively, eventually leading to model collapse. Building on this theoretical framework, we integrate our proposed token editing method and analyze whether our method can prevent model collapse.
Notation and Preliminaries
For a given distribution $P_{\Sigma,w,\sigma^{2}}$, the data $(x,y)\sim P_{\Sigma,w,\sigma^{2}}$ on $\mathbb{R}^{d}\times\mathbb{R}$, where $x$ is drawn from a multivariate normal distribution $x\sim\mathcal{N}(0,\Sigma)$, $\epsilon$ is an independent noise term sampled from $\mathcal{N}(0,\sigma^{2})$, and the label $y$ is given by the linear model $y=x\cdot w^{*}+\epsilon$.
Iterative Data Editing Process
We utilize the model obtained from the previous round of training to make a limited number of modifications. Specifically, we resample and replace data points with relatively high confidence. The editing operations are defined by the matrices $\{M_{1},M_{2},\dots,M_{n}\}$. The iterative data synthesis and model-fitting process is formalized as follows:
@@
Model collapse is ultimately reflected through test error. Following previous work, we adopt the standard test error definition as presented in (Gerstgrasser et al., 2024). For any linear estimator $\hat{w}$ derived from the training data, we evaluate the test error using the standard method as follows:
$E_{test}(w)\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}\mathbb{E}\left[(x_{test}^{T}w-y_{test})^{2}\right]-\sigma^{2}=\mathbb{E}[\|w-w^{*}\|_{\Sigma}^{2}]$ (4)
@@ 3.3 Test Error Under Data Editing
Theorem 2
$E_{test}(\hat{w}_{n+1})\leq\frac{2\sigma^{2}d}{T-d-1}$ (6)
Recalling that the cause of model collapse (Dohmatob et al., 2024a): training iteratively on synthetic data leads to an accumulation of error over iterations, as shown in the following equation:
Comparing Eq. 6 with Eq. 8, the test error under data editing is bounded by a fixed value, preventing continuous error accumulation and thus avoiding model collapse. Based on the theoretical derivations and statistical analysis of synthetic data (§ 2.1), the underlying reason is that our approach retains the coverage of the initial distribution. We move away from pure data synthesis toward token-level editing, which allows us to obtain better data while avoiding model collapse. Moreover, remarkable previous studies (Dohmatob et al., 2024c; Gerstgrasser et al., 2024) pointed out similar conclusions. They indicated mixing real data with synthetic data will break model collapse and provide an upper bound under data accumulation. Different from their work, our data editing aims to yield better data, enabling synthetic data to perform well both in theory and practice, not only avoiding model collapse.
4 Experiments
To validate our proposed method, we conduct experiments across three stages of language model training including: pre-training, continual pre-training (CPT) and supervised fine-tuning (SFT).
4.1 Implementation
We use the Llama-3-8B (AI@Meta, 2024) as a prior distribution to estimate the token distribution in each text sample. The modification probability is set to $p=0.99$. This means that we resample tokens in positions where the probability exceeds $p$, and the resampling is based on the conditional probability given the preceding context. The entire process requires only a single forward pass, without auto-regressive generation. We integrate the fast inference engine vLLM (Kwon et al., 2023), allowing the entire data editing process to be completed on a single 4090 GPU. After completing the data editing, we compare the original data and the edited data on language model training performance across pre-training, CPT, and SFT. Here, we used top-k as the sampling strategy with $k=8$. We also experimented with top-p and rejection sampling, which produced similar results.
4.2 Datasets and Models
We provide an overview of our experimental setup. More training details are presented in Appendix C. As for pre-training, we pre-train the 1B OLMo model (Groeneveld et al., 2024) from scratch, using Dolma-sampled V6 (6B tokens) as the pre-training corpus. We use 8 general tasks in lm-evaluation-harness (Gao et al., 2024) to evaluate for pre-training models. As for continual pre-training, we follow Cheng et al. (2024a; b; c) to continual pre-train the OLMo-1B (Groeneveld et al., 2024) and Llama-3-8B (AI@Meta, 2024) on Biomedicine, Finance and Math. Each domain corpus contains 1B tokens. Correspondingly, we evaluate the continual pre-training models using 15 downstream tasks, with 5 tasks from each domain. As for supervised fine-tuning, we fine-tune Llama-3-8B on instruction tuning and code reasoning tasks, including 6 datasets in total. We evaluate the SFT models using 9 downstream tasks designed to measure instruction-following capabilities. All Llama-3-8B experiments use LoRA (Hu et al., 2021), while the OLMo-1B model is trained with full parameters.
@@ 4.2 Datasets and Models
Models MQP ChemProt PubMedQA RCT USMLE Average
@@
Models HeadLine FPB FiQA-SA ConvFinQA NER Average
@@
Models ARC-c GPQA GSM8K MATH MMLU Average
@@
4.3 Results
Table 2,  3, and  4 respectively demonstrate the effectiveness of our method in continual pre-training, pre-training, and fine-tuning tasks. Across these three stages of language model training, our method enhances the model’s performance on downstream tasks without increasing the data size. Our method unlocks the potential of potential of existing data, demonstrating that semi-synthetic data is a viable path to improving model performance.
Specifically, as shown in Table 2, our method shows consistent improvements over the source data across OLMo-1B and LLaMA-3-8B. For instance, in the Biomedicine domain, the average score for OLMo-1B increased from 36.63 to 40.89 with ToEdit, while LLaMA-3-8B saw an increase from 54.13 to 56.48. Table 3 further supports the effectiveness of our approach in pre-training. The average performance of OLMo-1B increases from 32.75 to 33.11, reflecting improved generalization capabilities. While the improvement is modest, the consistent trend across tasks like PIQA, BoolQ, and ARC-c highlights the broader applicability of our method.
As for SFT results in Table 4, using both the original and edited data, the results indicate a small but consistent improvement. Specifically, ToEdit improves original FLAN v2, with average performance increasing from 70.18 to 70.65. For Natural Instructions, the average performance of LLaMA-3-8B improves from 69.34 to 69.70, with gains in tasks like Winogrande and SIQA. These improvements demonstrate the adaptability of our method to instruction-tuning tasks. For code-related tasks, the improvements are particularly evident in ARC-c and GPQA, indicating better reasoning and code comprehension.
Table 3: General performance of the pre-trained base models. PT indicates we pre-train OLMo-1B from scratch. Experimental results demonstrate that our method can also enhance the effectiveness of pre-training.
PIQA BoolQ OBQA ARC-c ARC-e HellaSwag SIQA Winogrande Average
OLMo-1B (PT) 53.97 38.26 12.20 17.23 28.36 26.02 34.80 51.14 32.75
$\Delta$ ToEdit 54.13 38.65 12.80 18.43 27.48 25.94 34.95 52.49 33.11
Table 4: Performance of the SFT models. We fine-tune LLaMA-3-8B using instruction tuning and code reasoning tasks, comparing performance with the edited version produced by our method. The experimental results indicate that our approach can enhance the data for instruction-tuning and code reasoning tasks.
Models PIQA BoolQ HellaSwag SIQA Winogrande Average
@@ 4.3 Results
Natural Instructions Llama-3-8B 79.82 87.06 58.32 46.83 74.66 69.34
$\Delta$ ToEdit 80.58 87.80 58.27 46.93 74.90 69.70
CoT Llama-3-8B 79.87 81.28 59.72 49.69 74.51 69.01
@@
Open Assistant 1 Llama-3-8B 79.65 83.18 60.51 48.52 74.11 69.19
$\Delta$ ToEdit 79.98 83.91 60.34 48.31 74.66 69.44
Models ARC-c GPQA GSM8K MMLU Average
@@
OSS-Instruct-75K Llama-3-8B 51.28 27.46 49.58 62.14 45.76
$\Delta$ ToEdit 51.79 28.79 49.36 62.04 46.13
Evol-Instruct-110K Llama-3-8B 52.90 27.90 50.87 62.40 46.62
$\Delta$ ToEdit 52.22 29.69 50.87 62.60 46.92
@@ 4.4 Ablation Studies
We conduct 4 experiments on hyper-parameter $p$, including: (1) ablation studies on different values, (2) token percentage statistics, (3) comparisons of sampling strategies, and (4) an ablation study on sampling size.
As Table 7 shows different values of $p$ influences on BioMed, with fluctuations observed across various settings. The Table 8 presents the distribution percentages across different probability value ranges. As mentioned above, we need to refine the data while preserving mainly source distribution.
Table 5: Results of different sampling strategies.
Sampling Strategy PubMedQA MedMCQA MedQA (4 options)
Top-k 64.5 26.13 24.82
Top-p 63.8 27.11 25.61
Reject Sampling 64.5 28.90 28.20
As shown in Figure 6, a larger $p$ indicates fewer tokens will be resampled, while a smaller $p$ results in more tokens being resampled. Balancing performance and the preservation of data distribution, we set $p=0.99$ as threshold for our experiments.
Table 6: Ablation study on sampling size $k$ for top-k.
Sampling Size ($k$) PubMedQA MedMCQA MedQA (4 options)
$k=8$ 64.5 26.13 24.82
$k=64$ 63.8 28.14 27.34
Table 5 shows the results of different sampling strategies. Specifically, to control variables, we set $k=8$ for top-k sampling and $p=0.99$ for top-p sampling. We use reject sampling implementation in Liu et al. (2023). The results of reject sampling, top-p, and top-k are comparable. However, top-p involves a dynamic sampling range, and reject sampling requires multiple rounds of computation, leading to increased overhead. Considering computational efficiency, we chose top-k for sampling. This aligns with our original objective of maintaining minimal computational overhead. This aligns with our initial objective of minimizing computational overhead as much as possible. The Table 6 shows the ablation study on sampling size of top-k. The improvement achieved with larger values is relatively small. Therefore, we chose $k=8$ in our experiments.
Table 7: Performance impact of different resampled token condition ($p$) in Biomedicine domain.
PubMedQA MQP RCT USMLE ChemProt Avg
$p\geq 0.99$ 64.5 55.73 30.95 27.65 14.6 38.69
$p\geq 0.999$ 63.6 55.4 29.09 28.12 16.2 38.48
$p\leq 0.1$ 62.4 51.47 25.6 29.14 10.0 35.72
$p\leq 0.01$ 65.4 54.91 28.19 27.80 11.0 37.46
$p\leq 0.001$ 64.2 56.39 35.0 27.80 12.4 39.16
Table 8: Token distribution across different probability ranges in BioMed dataset.
Probability Range Percentage Token Count
0.0-0.1 34.7% 388,626,330
0.1-0.2 8.1% 90,716,809
0.2-0.3 5.4% 60,477,872
0.3-0.4 4.4% 49,278,266
0.4-0.5 3.8% 42,558,503
0.5-0.6 3.6% 40,318,546
0.6-0.7 3.7% 41,438,924
0.7-0.8 4.0% 44,798,424
0.8-0.9 5.2% 58,238,944
0.9-1.0 27.1% 303,543,988
5 Discussion
5.1 What Is the Difference Between Non-Iterative and Iterative Model Collapse?
From a setting perspective: The difference between the two lies in their scope. Non-iterative model collapse is not confined to training on self-generated data, allowing it to uncover broader properties of synthetic data. For instance, in our experiments, we train GPT-2 on the Cosmopedia dataset in a single generation, which was generated by Mixtral-8x7B-Instruct-v0.1. In contrast, iterative model collapse focuses on training the model over multiple generations using self-generated data.
From a property perspective: The non-iterative model collapse emphasizes the gap between human data and general purely synthetic data, particularly regarding distributional properties and n-gram features. In contrast, the iterative model collapse illustrates the iterative evolution of the model, resembling a self-play process. This process illustrates the gradual evolution of self-generated data. It does not involve an analysis of the differences in nature between self-generated and human-produced data. They both ultimately lead to model collapse, driven by the same underlying cause—synthetic data, though they investigate different aspects of synthetic data. The most common setting is training a model on a mixture of human and synthetic data, where the synthetic data is not generated by the model itself, and its exact origin may be unknown. Moreover, there are already numerous popular datasets, such as UltraChat and OpenOrca, that combine synthetic and human-produced data to improve training diversity and robustness.
5.2 Definition and Characteristics of Synthetic Data
5.3 Why Does the Observed Probability Distribution Exhibit Filtering Potential?
From the perspective of information theory, we can analyze the filtering potential of the U-shape distribution as follows: We utilize the U-shape distribution in Figure 6 to re-sample tokens in the high-probability region, to adjust the U-shaped distribution toward a uniform distribution. By doing so, we can maximize the information entropy. According to information theory, maximizing information entropy is achieved when the distribution is uniform. Lemma 1: Let $X$ be a discrete random variable with $n$ possible outcomes. If the probability of each outcome is uniform, i.e., $P(x_{i})=\frac{1}{n}$ for all $i\in\{1,2,\dots,n\}$, the Shannon entropy is maximized, given by:
$H(X)$ $=-\sum_{i=1}^{n}\frac{1}{n}\log\frac{1}{n}=\log n.$ (10)
5.4 Gradual Decline in Editing
Table 9: Percentage of tokens requiring edits in the Natural-Instructions dataset. The total number of tokens is 4,671,834. and “Gen” is short for “Generation”.
Gen 1 (source) Gen 2 Gen 3
Tokens ($p>0.99$) 584,103 549,519 517,433
Percentage 12.5% 11.76% 11.08%
We present the percentage statistics of edited tokens in Table 9, demonstrating that the edited tokens indeed exhibit a progressive decrease. Specifically, We observe that the percentage of edited tokens (above the threshold $p>0.99$) decreases as the generation number increases. Theoretically, this is a process of distribution shifting. When tokens ($p>0.99$) are resampled, randomness is introduced. The sampling process can select tokens with lower probabilities. Then, tokens ($p>0.99$) is replaced, leading to a reduction of edited tokens in subsequent generations. The Table 9 provides empirical evidence for this pattern of decay.
6 Related Work
Model Collapse
Shumailov et al. (2024); Dohmatob et al. (2024a; c) demonstrate AI models trained recursively on data generated by earlier versions of themselves over time can result in performance degradation, ultimately rendering the AI model completely useless. This process can be formulated as follows:
$E_{test}(\hat{w}_{n+1})=\frac{\sigma^{2}d}{T-d-1}\times n$
This indicates that the error will continuously increase with the number of iterations $n$. Dohmatob et al. (2024c) further point out that synthetic data also contribute to a truncation of the scaling law. This phenomenon stems from the sampling strategy (e.g., Top-p) used during the language model’s generation process. Gerstgrasser et al. (2024) further adjust the data iteration setting by replacing data replacement with data accumulation during the iterative process. They demonstrate that data accumulation can prevent model collapse. Inspired by the above work, we believe that training language models on synthetic datasets will be inevitable in the future. Therefore, it is crucial to theoretically discuss how to prevent model collapse. Building on the above theoretical framework, we prove that token-level editing establishes an upper bound during the iterative process, thereby preventing the continuous accumulation of errors.
Synthetic Data
Phi-1/2 (Gunasekar et al., 2023) demonstrate that the synthetic data can boost training efficiency and performance compared to raw data in language model pre-training. Liu et al. (2024) highlight that synthetic data will play a crucial role in the development of AI. For example, synthetic data can be used to construct highly specialized datasets, enhancing the performance of downstream tasks. Trinh et al. (2024) utilize synthetic math data to train a 125M language model, which successfully solved 25 out of 30 selected problems from the International Mathematical Olympiad (IMO) problem set. Zhang et al. (2024) develop a biomedical instruction dataset that was used to train specialized bio-models, enabling them to excel in answering questions related to medical exams and clinical scenarios. Eldan & Li (2023) introduce a novel synthetic dataset and evaluation paradigm that enables small language models to generate coherent, diverse, and grammatically sound stories. As outlined above, in the post-training stages of LLMs, synthetic data enhances downstream task performance and aligns foundation models with humans. And Maini et al. (2024) propose rephrasing the pre-training data into a Wikipedia or Q/A style to achieve better alignment with downstream tasks. Synthetic data is a powerful tool for training. Our approach is also based on synthetic data methods. Instead of sampling data solely based on this prior, we modify the data using the prior as a guide.

### v2

@@ 4 Token-Level Editing
We introduce token-level editing as a method for generating semi-synthetic data. Furthermore, we present a theoretical analysis and proof demonstrating that the test squared error of our method has a finite upper bound, regardless of the number of iterations. This ensures prevention of model collapse while enhancing performance.
@@ 4.1 Method
$P(y_{1},\dots,y_{n}\mid\textbf{x})=\prod_{i=1}^{n}P(y_{i}\mid\dots,y_{<i},\textbf{x}).$ (1)
@@
Inspired by previous studies of data selection (Mindermann et al., 2022; Ankner et al., 2024; Lin et al., 2024), prior distributions can serve as pointers indicating useless or learnable samples. In this case, we use a pre-trained language model to infer the pre-training corpus. As illustrated in Figure 5, even a model pre-trained on trillions of tokens can not fit the pre-training corpus perfectly. Specifically, $75\%$ is under 0.6 probability. The tokens with both high and low probabilities are the most concentrated, suggesting the potential for data filtering. We leverage this U-shape distribution as a pointer to resample tokens. Specifically, we use a language model as prior distribution to compute each token’s conditional probability $P(\cdot|\textbf{x})$ if the probability exceeds a certain threshold $P(\cdot|\textbf{x})\geq p$, it indicates that the token is easy to learn, and we proceed with resampling at that point. The filtering potential of the U-shaped distribution is discussed in § G.2.
@@ 4.2 Theoretical Analysis
To gain deeper mathematical insights, we utilize an analytical framework of the linear model and adopt notations in prior research (Mobahi et al., 2020; Dohmatob et al., 2024a; Gerstgrasser et al., 2024). This theoretical framework primarily considers a linear model that iteratively trains on its own generated data, similar to pipelines like self-play and self-distillation, but without complex constraints. The process involves training continuously on the data generated by the previous generation of the model. (Dohmatob et al., 2024a) point out that with iterative training, test errors accumulate progressively, eventually leading to model collapse. Building on this theoretical framework, we integrate our proposed token editing method and analyze whether our method can prevent model collapse.
Notation and Preliminaries For a given distribution $P_{\Sigma,w,\sigma^{2}}$, the data $(x,y)\sim P_{\Sigma,w,\sigma^{2}}$ on $\mathbb{R}^{d}\times\mathbb{R}$, where $x$ is drawn from a multivariate normal distribution $x\sim\mathcal{N}(0,\Sigma)$, $\epsilon$ is an independent noise term sampled from $\mathcal{N}(0,\sigma^{2})$, and the label $y$ is given by the linear model $y=x\cdot w^{*}+\epsilon$.
Iterative Data Editing Process We utilize the model obtained from the previous round of training to make a limited number of modifications. Specifically, we resample and replace data points with relatively high confidence. The editing operations are defined by the matrices $\{M_{1},M_{2},\dots,M_{n}\}$. The iterative data synthesis and model-fitting process is formalized as follows:
@@
Model collapse is ultimately reflected through test error. Following previous work, we adopt the standard test error definition as presented in (Gerstgrasser et al., 2024). For any linear estimator $\hat{w}$ derived from the training data, we evaluate the test error using the standard method:
$E_{test}(w)\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}\mathbb{E}\left[(x_{test}^{T}w-y_{test})^{2}\right]-\sigma^{2}=\mathbb{E}[\|w-w^{*}\|_{\Sigma}^{2}],$ (4)
@@ 4.3 Test Error Under Data Editing
Table 2: General performance of the pre-trained base models. PT indicates we pre-train OLMo-1B from scratch.
PIQA BoolQ OBQA ARC-c ARC-e HellaSwag SIQA Winogrande Avg
OLMo-1B (PT) 53.97 38.26 12.20 17.23 28.36 26.02 34.80 51.14 32.75
$\Delta$ ToEdit 54.13 38.65 12.80 18.43 27.48 25.94 34.95 52.49 33.11
@@
Models MQP ChemProt PubMedQA RCT USMLE Avg
@@
Models HeadLine FPB FiQA-SA ConvFinQA NER Avg
@@
Models ARC-c GPQA GSM8K MATH MMLU Avg
@@
Theorem 2.
$E_{test}(\hat{w}_{n+1})\leq\frac{2\sigma^{2}d}{T-d-1}$ (6)
$E_{test}(\hat{w}_{n+1})$ $\leq\frac{\sigma^{2}d}{T-d-1}$
$\quad+\sigma^{2}\sqrt{\mathbb{E}\left[\text{tr}\left((X^{\top}X)^{-2}\right)\right]}\cdot\frac{\sqrt{\mathbb{E}\left[\text{tr}(M_{1})\right]}}{1-\eta}.$
We provide supporting evidence for the assumption in § G.3. Recalling model collapse (Dohmatob et al., 2024a): training iteratively on synthetic data leads to an accumulation of error over iterations, as shown in the following equation:
Comparing Eq. 6 with Eq. 7, the test error under data editing is bounded by a fixed value, preventing continuous error accumulation and thus avoiding model collapse. Based on the theoretical derivations and statistical analysis of synthetic data (§ 3.1), the underlying reason is that our approach retains the coverage of the initial distribution. We move away from pure data synthesis toward token-level editing, which allows us to obtain better data while avoiding model collapse. Moreover, remarkable previous studies (Dohmatob et al., 2024c; Gerstgrasser et al., 2024) pointed out similar conclusions. They indicated mixing real data with synthetic data will break model collapse and provide an upper bound under data accumulation. Different from their work, our data editing aims to yield better data, enabling synthetic data to perform well both in theory and practice.
Table 4: Performance of the SFT tasks. We fine-tune LLaMA-3-8B using instruction tuning and code reasoning tasks, comparing performance with the edited version produced by our method. HS and WG are short for HellaSwag and Winogrande respectively.
Models PIQA BoolQ HS SIQA WG Avg
@@
Natural Llama-3 79.82 87.06 58.32 46.83 74.66 69.34
Ins. $\Delta$ ToEdit 80.58 87.80 58.27 46.93 74.90 69.70
CoT Llama-3 79.87 81.28 59.72 49.69 74.51 69.01
@@
Open Llama-3 79.65 83.18 60.51 48.52 74.11 69.19
Assist. $\Delta$ ToEdit 79.98 83.91 60.34 48.31 74.66 69.44
Models ARC-c GPQA GSM8K MMLU Avg
@@
OSS- Llama-3 51.28 27.46 49.58 62.14 45.76
Inst. $\Delta$ ToEdit 51.79 28.79 49.36 62.04 46.13
Evol- Llama-3 52.90 27.90 50.87 62.40 46.62
Ins. $\Delta$ ToEdit 52.22 29.69 50.87 62.60 46.92
5 Experiments
We validate our method in three language model training stages: pre-training from scratch (PT), continual pre-training (CPT), and supervised fine-tuning (SFT).
5.1 Implementation
We use the Llama-3-8B (AI@Meta, 2024) as a prior distribution to estimate the token distribution in each text sample. The modification probability is set to $p=0.99$. This means that we resample tokens in positions where the probability exceeds $p$, and the resampling is based on the conditional probability given the preceding context. The entire process requires only a single forward pass, without auto-regressive generation. We integrate the fast inference engine vLLM (Kwon et al., 2023), allowing the entire data editing process to be completed on a single 4090 GPU. We use top-k as the sampling strategy with $k=8$. We also incorporate top-p sampling and rejection sampling in our ablation studies.
5.2 Datasets and Models
We provide an overview of our experimental setup, more details are in Appendix F. For pre-training, we pre-train the 1B OLMo model (Groeneveld et al., 2024) from scratch using Dolma-sampled V6 (6B tokens) and evaluate on 8 general tasks. For continual pre-training, we follow (Cheng et al., 2024a, b, c) to continual pre-train OLMo-1B (Groeneveld et al., 2024) and Llama-3-8B (AI@Meta, 2024) on corpora of Biomedicine, Finance, and Math, evaluating on 5 downstream tasks per domain. For supervised fine-tuning, we fine-tune Llama-3-8B on instruction tuning and code reasoning tasks, evaluating on 9 downstream tasks.
5.3 Main Results
Table 23, and 4 respectively demonstrate the effectiveness of our method in pre-training from scratch, continual pre-training and fine-tuning tasks. Across these three stages of language model training, our method consistently enhances the model performance on downstream tasks without increasing the data size. This consistency is validated across two models. This indicates that our method unlocks the potential of existing data, demonstrating that semi-synthetic data is a viable path to improving model performance. A further numerical analysis is provided in § C.
Table 5: Ablations on resampled token condition ($p$) in biomedicine domain.
PubMedQA MQP RCT USMLE ChemProt Avg
$p\geq 0.99$ 64.50 55.73 30.95 27.65 14.60 38.69
$p\geq 0.999$ 63.60 55.40 29.09 28.12 16.20 38.48
$p\leq 0.1$ 62.40 51.47 25.60 29.14 10.00 35.72
$p\leq 0.01$ 65.40 54.91 28.19 27.80 11.00 37.46
Table 6: Token distribution across different probability intervals in the biomedicine domain dataset.
Interval Percent # Tokens Interval Percent # Tokens
$[0.0,0.1)$ 34.7% 389M $[0.5,0.6)$ 3.6% 40M
$[0.1,0.2)$ 8.1% 91M $[0.6,0.7)$ 3.7% 41M
$[0.2,0.3)$ 5.4% 60M $[0.7,0.8)$ 4.0% 44M
$[0.3,0.4)$ 4.4% 49M $[0.8,0.9)$ 5.2% 58M
$[0.4,0.5)$ 3.8% 43M $[0.9,1.0)$ 27.1% 303M
Table 7: Ablations on sampling strategy.
Strategy PubMedQA MedMCQA MedQA
Top-k 64.50 26.13 24.82
Top-p 63.80 27.11 25.61
Rejection Sampling 64.50 28.90 28.20
Table 8: Ablations on sampling size $k$ for top-k.
Sampling Size ($k$) PubMedQA MedMCQA MedQA
$k=8$ 64.50 26.13 24.82
$k=64$ 63.80 28.14 27.34
@@ 5.4 Ablation Studies
We conduct experiments on hyper-parameter $p$, including: (1) ablation studies on different values, (2) token percentage statistics, (3) comparisons of sampling strategies, and (4) an ablation study on sampling size.
Table 5 shows the impact of different $p$ values on the model performance, with fluctuations observed across various settings. Table 6 presents the distribution percentages across different probability value ranges. As mentioned above, we need to refine the data while preserving mainly source distribution. As shown in Figure 5, a larger $p$ indicates fewer tokens will be resampled, while a smaller $p$ results in more tokens being resampled. To balance model performance and data distribution preservation, we set $p=0.99$ as threshold for our experiments. Table 7 presents the results of different sampling strategies. Specifically, to control variables, we set $k=8$ for top-k sampling and $p=0.99$ for top-p sampling. We use rejection sampling implementation in (Liu et al., 2023). The results of reject sampling, top-p, and top-k are comparable. However, top-p involves a dynamic sampling range, and reject sampling requires multiple rounds of computation, leading to increased overhead. Considering computational efficiency, we choose top-k for sampling. Table 8 shows the ablation study on sampling size of top-k. The performance gain from increasing $k$ is relatively small. Therefore, we set $k=8$ in our experiments. And a detailed case for token editing is provided in Table 12.